OpenStudy (anonymous):

I want to solve an equation for wavelength, but my algebra is a little rusty. Given the following equation:\[\tilde{\nu}=\frac{1}{\lambda}=R\bigg(\frac{Z^2}{n^2_i}-\frac{Z^2}{n^2_i}\bigg)\] Is there a better way to solve for λ than this? \[\lambda=\frac{1}{\nu}=\frac{1}{R(Z^2/n^2_i-Z^2/n^2_f)}\]

OpenStudy (anonymous):

It seems like I should be able to have everything on the either the top or bottom of the fraction, instead of embedding fractions.

OpenStudy (anonymous):

I plugged this problem into WolframAlpha, and it gave me an idea for simplifying it. Does this work?\[\lambda = \frac{1}{\tilde \nu}=-\frac{n_i^2n_f^2}{RZ^2(n_i^2-n_f^2)} = -\bigg(\frac{1}{R(n_i^2-n_f^2)}\bigg)\bigg(\frac{n_in_f}{Z}\bigg)^2\]

OpenStudy (anonymous):

The above is correct, however, I talked to the people in the math OpenStudy group, and evidently this is the most compact way to write the formula to solve for the wavelength in the Bohr model: \[λ=\tilde \nu^{−1}=(RZ^2(n^{−2}_i−n^{−2}_f))^{−1}\]